OpenStudy (amy0799):

if x^2+y^2=a where a is a non-zero constant, which of the following conditions are necessary for

OpenStudy (amy0799):

\[\frac{ d ^{2}y }{ dx ^{2} }>0\] a. y<0 b. y>0 c. x>0

zepdrix (zepdrix):

Have you tried to find your y''? :)

OpenStudy (amy0799):

yes

OpenStudy (amy0799):

\[y''=-\frac{ x ^{2}+y ^{2} }{ y ^{3}}\]

zepdrix (zepdrix):

Ooo ok ya :O derivative looks good! So we have to figure out some conditions huh? Hmm

OpenStudy (amy0799):

yes I don't know how to figure that out

zepdrix (zepdrix):

Notice that x^2 is ALWAYS positive. So y is the only thing dictating the sign of y''.

zepdrix (zepdrix):

And further than that, y^2 doesn't affect our sign either. So we need only to pay attention to the negative in front, and the y^3

zepdrix (zepdrix):

\[\large\rm y''=-\frac{c}{y^3}\] Where c is a positive number. y'' is positive when \(\large\rm -\frac{c}{y^3}\gt0\)

zepdrix (zepdrix):

So ummmm

zepdrix (zepdrix):

When you cube a negative, you get an extra negative sign popping out, ya? So I guess we just need to make sure that y is negative, in order for the two negatives to cancel out. Make sense? :o

zepdrix (zepdrix):

There's probably a more algebraic way to do this, but I can't think of it :) lol

OpenStudy (amy0799):

so b is the answer if I'm understanding correctly?

OpenStudy (amy0799):

wait no, its A

zepdrix (zepdrix):

It's certainly not c, the sign of x has no effect on our second derivative. A? When y is less than 0? So y has to be negative for the entire expression to be positive? Yay good job \c:/

OpenStudy (amy0799):

so a is the answer?

zepdrix (zepdrix):

Here is a way you can visualize it. Recall that the second derivative tells us about concavity. So a positive second derivative should represent a shape of `concave up`. When y is negative, we're dealing with the BOTTOM HALF OF THE CIRCLE. The shape of which is concave up! :) ya?

zepdrix (zepdrix):

yes

OpenStudy (amy0799):

ooh ok. Thank you for the great explanation!

zepdrix (zepdrix):

np c: